OpenStudy (anonymous):

Find the area of the largest rectangle with a base on the positive x axis, its right side on the line x=9 and which is inscribed under the curve f(x) = root(x)

OpenStudy (anonymous):

I don't get it... can anybody help please?

OpenStudy (mattfeury):

you have to integrate on root(x) from x=0 to x=9.

OpenStudy (mattfeury):

actually that doesn't give you a rectangle. hmmm....

OpenStudy (anonymous):

elaborate please

OpenStudy (mattfeury):

this is the graph. it is strange though because at x=0, y = 0

OpenStudy (anonymous):

screen shot is nothing

OpenStudy (anonymous):

hmm, i can't see anything

OpenStudy (mattfeury):

*hopefully*

OpenStudy (mattfeury):

you have to find the largest rectangle possible. looking at the graph i would guess (4,2) but that's just a though....

OpenStudy (anonymous):

apparently we have to do it through optimization - derivatives and all that fun stuff

OpenStudy (anonymous):

This is one of those questions you ask an instructor whose got over 20 yrs cal experience.

OpenStudy (mattfeury):

find a function for the area of the square and optimize. the area of the square at a given x up until 9 = x*y = sqrt(x) * (9-x)

OpenStudy (mattfeury):

try deriving that and see if find an extrema

OpenStudy (anonymous):

not really :) as long as you have the skill then it's no prob

OpenStudy (anonymous):

it's one of those questions that are a challenge ^_^

OpenStudy (anonymous):

loving the enthusiasm sstarica, but i'm getting mercilessly owned in cal a

OpenStudy (anonymous):

LOL, nah, let it be the opposite. It's quite simple, believe me :)

OpenStudy (anonymous):

logically, I'd say the following: since the lenght of the rectangle is = 9, and is under the sqrt(x) now, let's first draw sqrt of (x) and notice that when you take x = 9, you'llhave y = 3. Logically, the rectangle won't exceed this limit, so the width is equal = 3

OpenStudy (anonymous):

length*

myininaya (myininaya):

matt has ir right! good job

OpenStudy (anonymous):

since it's inscribed under it

OpenStudy (mattfeury):

yes the maxima of that graph is 3!

OpenStudy (anonymous):

:) no need for all this.

myininaya (myininaya):

now we need to find A'

OpenStudy (anonymous):

all you had to do is draw sqrt(x) and see the limit of the rectangle. ^_^

OpenStudy (anonymous):

why A'? doesn't he want A?

OpenStudy (anonymous):

A = 3 x 9 = 27 that if he asked for A right? ^_^

myininaya (myininaya):

to maximize you find A'

OpenStudy (mattfeury):

so it's 3 * sqrt(3)

OpenStudy (mattfeury):

or no. 6 * sqrt(3)

OpenStudy (anonymous):

oh right lol my bad

myininaya (myininaya):

yes x=3 is right matt thats are only critcal number thats in the domain of A

OpenStudy (anonymous):

wait wait, why is 9=xy? doesn't A=xy?

OpenStudy (anonymous):

since he wants the largest area, then yes maximize, proceed matt and hope you understood what was going on Ire :)

myininaya (myininaya):

i didnt see A=9 up there no

OpenStudy (anonymous):

you want the largest area, so you have to maximize in this case, they didn't ask for the "area" alone, but largest one

OpenStudy (mattfeury):

right. to find where the area is a maximum, take a function for the area and find the maxima point. take that point to make your rectangle. since you go up to 9 the x width is (9-3) = 6. the height of the rectangle is sqrt(3)

OpenStudy (anonymous):

could you remind me how to find the maxima point please?

OpenStudy (anonymous):

which means "optimize" lol, alight matt will take it from here ^_^ good luck

OpenStudy (mattfeury):

take the derivative of the area function 'A'. find where the derivative = 0.

OpenStudy (mattfeury):

normally, you'd have to do the second derivative test to find out if it is a max/min too. but i know secrets.

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

1. A = (9-x) * sqrt(x)

OpenStudy (anonymous):

wow, thanks guys

OpenStudy (mattfeury):

very nice.

myininaya (myininaya):

also i never use the second derivative test it is pointless you can use A' to see if it is increasing to decreasing at x=3 to find that it is a max

myininaya (myininaya):

The first derivative rocks!

OpenStudy (anonymous):

thanks for the help guys, really appreciated